OpenStudy (anonymous):

Simplify f(x) = ((4x^2)+7x-2)/(4x-1) Using complete sentences, explain why f(1) = 3, f(0) = 2, and f(-1) = 1, yet f(1/4) is undefined. Be sure to show your work.

OpenStudy (anonymous):

can you factorize the part 4x^2 + 7x - 2 ?

OpenStudy (anonymous):

I'm not sure how to

OpenStudy (anonymous):

find 2 numbers, if multiplied, the result will be 4*(-2) = -8, and if added, the result will be 7 

OpenStudy (anonymous):

8 and -1?

OpenStudy (anonymous):

correct. next, can you find a and b from this? 

OpenStudy (anonymous):

(4x+8)(x-1) 8*-1=-1?

OpenStudy (anonymous):

not the factors yet can you find a and b from the drawing?

OpenStudy (anonymous):

a*(-1) = -1 4*b=8

OpenStudy (anonymous):

a=1 b=2